OpenStudy (anonymous):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

hero (hero):

1. First find the slope of the other line by substituting each point separately into y = mx + b: -11 = -4m + b 13 = 2m + b 2. Isolate b in each equation: b = 4m - 11 b = 13 - 2m 3. Set b = b 4m - 11 = 13 - 2m 4. Solve for m: 4m + 2m = 13 + 11 6m = 24 m = 24/6 m = 4 Since m represents slope, then the slope of the line is 4. Since the slope of the other line is parallel and parallel lines have the same slope, then the slope of the other line is also 4.